Write the formula for the compound that will form when the following pairs of ions are mixed then classify the compound as a type I or type II.
Mg2+, Cl1-
Ti4+, Cl1-
Ag1+, Br1-
Ba2+, NO31-
Fe2+, HO1-
Ni4+, S2-
Zn2+,SO42-
Sr2+, P3-
Pb2+, N3-
Al3+, Cl1-
1
Expert's answer
2020-07-24T14:52:53-0400

Mg2+, Cl1- = MgCl2

Ti4+, Cl1- = TiCl4

Ag1+, Br1- = AgBr

Ba2+, NO31- = Ba (NO3)2

Fe2+, HO1- = Fe(OH)2

Ni4+, S2- = (NH4)2S .

Zn2+,SO42- = Zn ( SO4)

Sr2+, P3- = Sr3(P)2

Pb2+, N3- = Pb3 (N)2

Al3+, Cl1- = Al (Cl)3